A ring of radius R has charge -Q distributed uniformly over it.A charge q is placed at the centre of the ring such that electric field becomes zero at a point on the axis of the ring at a distance R from centre of the ring. The value of q=Q/4rootW . Find W.

Grade:11

1 Answers

harsha vardhan
11 years ago

consider the sphere to have uniform charge distribution lambda

the field due to a small piece of the ring at a point at a distance r on the axial line is given by

dE= K* (lambda)dx/2*r*r

integrating

E=KQ/2*r*r

net field=0

so 

kq/r*r

q=Q/2

w=1/4